if vector a,b,c are unit vectors such that vector a is perpedicular to the plane of vector b ,c and the angle between vector b & vector c is pi/3 then |a+b+c| = ?

Grade:11

1 Answers

Khimraj
5 years ago
let x = |a + b + c|
then
x2 = a2 + b2 + c2 + 2a.b + 2b.c + 2a.c
x2 = 1 + 1 + 1 + 2*1*1*cos90 + 2*1*1*cos90 + 2*1*1*cos60
x2 = 1 + 1 + 1 + 0 + 0 + 1 = 4
so x = 2
Hope it clears. If you like answer then please approve it.